# LSAST: Enhancing Cybersecurity through LLM-supported Static Application Security Testing

### v1

@@ IV-C Preparing Vulnerability Reports for Knowledge Retrieval Process
The process begins with summarizing the functionality of each code snippet within the vulnerability reports and storing these summaries in a vector database. This task utilizes an LLM, and specific instructions can be found in Listing 6. When scanning the target code (TC) for vulnerabilities, we also summarize its functionality using the LLM. Subsequently, we conduct a similarity search in our vector database to find vulnerable code snippets (VC) with similar functionalities. Once we have gathered multiple functionalities of vulnerable code, we correlate them back to the respective vulnerability reports using the sequence numbers stored in the vector database entries. This method allows us to extract and deliver the most relevant vulnerability reports to the LLM for analysis.
Given the weak correlation between code functionality and vulnerabilities, we have developed a new approach for conducting similarity searches.
@@
A stronger correlation often emerges when examining code abstraction in relation to vulnerabilities. Abstraction, also known as code minimization, involves simplifying source code to its fundamental components essential for correct execution by compilers or interpreters. To implement a similarity search based on code abstraction rather than functionality, we begin by extracting the code from vulnerability reports and then applying code minification techniques. This process can be achieved using code minifiers or leveraging LLM capabilities, as detailed in Listing 7. After minifying the snippets of vulnerable code, we store them in a vector database.
@@
When we now scan the target code (TC) for vulnerabilities, we first minimize the TC and then search for vulnerable code (VC) with similar minified representations. Subsequently, we aggregate multiple code abstractions and correlate them back to the original reports using sequence numbers, as previously explained. This focused approach ensures that we only gather vulnerability reports with structures highly resembling the target code, increasing the likelihood of identifying identical vulnerabilities that are present in both VC and TC. This method aligns closely with the methodology employed by conservative SAST scanners, which aim to identify vulnerable patterns within code.
@@ IV-D Knowledge Retrieval Methods
In the following figures you can observe the various methods available for constructing the knowledge retrieval process through the combination of the functionality approach, abstraction approach and conservative SAST scanner results.
TC:
Target Code
SR:
SAST-Scanner Result
VR:
Vulnerability Reports
FVR:
Functionality-similar Vulnerability Reports
CVR:
Code-Abstraction-similar Vulnerability Reports
k:
Value of top $k$ in the similarity search
IV-D1 Own-Knowledge
Figure 1: Own-Knowledge Approach RAW
@@ IV-D1 Own-Knowledge
This method is the most basic of all. Here we just simply query the LLM directly by including TC in the scanning prompt, without using the knowledge retrieval system or bearer results.
IV-D2 Own-Knowledge-LSAST (our new approach)
Figure 2: Own-Knowledge Approach LSAST
@@ IV-D2 Own-Knowledge-LSAST (our new approach)
This method enhances the Own-Knowledge Approach with the bearer scanner results. We first run a bearer scan and then provide the results to the LLM, with the target code to analyze.
IV-D3 Functionality
Figure 3: Functionality Approach RAW
@@ IV-D3 Functionality
Here we use the LLM to summarize the functionality of the TC. We use the functionality of the TC to make a similarity search inside the vector database that contains all the functionalities of the vulnerable code snippets. From the similarity search we gather the most similar VC snippets, that are then used to gather the correlated VR’s. The VR’s from the similarity search are then provided to the LLM with the target code to analyze.
IV-D4 Functionality-LSAST (our new approach)
Figure 4: Functionality Approach LSAST
@@ IV-D4 Functionality-LSAST (our new approach)
Same as the functionality approach, but we also provide bearer scanner results to the LLM.
IV-D5 Abstraction (our new approach)
Figure 5: Abstraction Approach RAW
@@ IV-D5 Abstraction (our new approach)
We use the LLM to abstract the TC. We use the abstraction of the TC to make a similarity search inside the vector database that contains all the code abstractions of the vulnerable code snippets. From the similarity search we gather the most similar VC snippets, that are then used to gather the correlated VR’s. The VR’s from the similarity search are then provided to the LLM with the target code to analyze.
IV-D6 Abstraction-LSAST (our new approach)
Figure 6: Abstraction Approach LSAST
@@ IV-D6 Abstraction-LSAST (our new approach)
Same as the abstraction approach, but we also provide bearer scanner results to the LLM.
IV-D7 Combined (our new approach)
Figure 7: Combined Approach RAW
@@ IV-D7 Combined (our new approach)
This approach combines Functionality and Pattern similarity search. k VR’s from the Functionality similarity search and k VR’s from the Pattern similarity search are gathered. These vulnerability reports are then provided to the LLM.
IV-D8 Combined-LAST (our new approach)
Figure 8: Combined Approach LSAST
@@ IV-D8 Combined-LAST (our new approach)
Inspired by [9], which also incorporates industrial solutions as guidance, we also provide bearer scanner results to the LLM, and the other settings remain the same as the combined approach.
@@ V Evaluation
Since previous work has already shown that LLM’s can be utilized for vulnerability scanning and it is very time-consuming to evaluate all approaches, we have decided to structure our approach in the most efficient way. First, for all evaluations, we are using Llama-3-70B, as it is a high-performing mid-sized model that can be hosted without needing very large cloud infrastructures or sending the code to third-party LLM providers. Also, we have first evaluated which of the RAG approaches is most efficient in a small-scale evaluation before attempting the big evaluation between the raw models and the best-performing RAG approach.
V-A Abstraction vs Functionality vs Combined
@@
TABLE I: Comparison of Abstraction, Functionality, and Combined Approaches
We started by evaluating the Abstraction Approach (4.4.5), Functionality Approach (4.4.3), and the Combined Approach (4.4.7) without feeding the bearer input into the prompt. We evaluated our approaches using four well-known open-source security testing projects:
  1. 1.
    DVWA (Damn Vulnerable Web Application) [25]: A PHP/MySQL web application intentionally designed with vulnerabilities to help security professionals test their tools and hone their vulnerability detection skills.
  2. 2.
    DVNA (Damn Vulnerable NodeJS Application) [2]: A Node.js web application that intentionally incorporates OWASP Top 10 Vulnerabilities to demonstrate how these vulnerabilities can be identified and mitigated.
  3. 3.
    OWASP Juice Shop [17]: A modern JavaScript-based application (using Node.js, Express, and Angular) intentionally built with vulnerabilities. It serves as a platform for security training and testing of security tools.
  4. 4.
    WebGoat [18]: A Java-based web application deliberately including vulnerabilities, designed for security professionals to explore and understand common Java-specific security issues.
The projects include different programming languages like Java, JavaScript, TypeScript, and PHP and are built to include a wide range of different vulnerabilities. Across all 4 projects, using our 3 scanner approaches, we have scanned a total of 10 files that include more than 15 different kinds of vulnerabilities.
For our evaluation metrics, we used standard measures in vulnerability detection:
@@
As ground truth, we have decided to use the Bearer results, challenging the 3 approaches to find as many vulnerabilities as the static scanner finds. This evaluation is to find out which of the 3 approaches covers most of the underlying vulnerabilities.
In Table 1, you can see that the Abstraction Approach has outperformed the Functionality Approach, but the Combined Approach has outperformed both the Abstraction Approach and the Functionality Approach. The 100% FP-Rate is achieved since all files include at least 1 vulnerability, and therefore a scan never had the opportunity to label a file as false positive.
Interpretation
As explained earlier, there is a higher correlation between code abstraction and underlying vulnerabilities compared to functionality and underlying vulnerabilities. This is because functionalities can differ greatly but still have the same vulnerabilities, and vice versa. Conservative scanners also scan for underlying code patterns or syntactic features. Therefore, we can see that our new Abstraction and Combined approaches work better than the Functionality approach, which was introduced in LLM4Vuln [23].
@@
TABLE II: Comparison of similarity search for Abstraction and Functionality Approaches
In Table 2, you can see the results of the Similarity Search Performance. The TP-Rate can be interpreted as how many of the vulnerabilities that we passed to the LLM were also found by Bearer. In other words, around 67% of the reports that we passed to the LLM, with the abstraction approach, were related to other vulnerabilities than the ones found by Bearer.
Interpretation
We distract the LLM from finding the vulnerabilities that were identified by Bearer, by passing large numbers of vulnerabilities that were not found by Bearer to the LLM.
@@ V-B RAW vs LSAST vs LSAST-Combined
We have evaluated that the combined approach performs the best out of the three knowledge retrieval approaches. Now we want to compare the combined approach with the approaches that do not use a knowledge retrieval system. For this, we have again scanned the open-source projects DVWA, DVNA, Juice Shop, and WebGoat, but this time using the combined LSAST-Approach (4.4.8) (with knowledge retrieval and bearer results), the raw LSAST-Approach (4.4.2) (without knowledge retrieval, but with bearer results), and the raw Approach (4.4.1) (without knowledge retrieval and without bearer results). Since we now pass the Bearer results directly to the LLM for the combined LSAST-Approach and the raw LSAST-Approach, we want to analyze how many true vulnerabilities the scanners can find besides the ones Bearer already finds. To determine if the found vulnerabilities are true positives or false positives, we have gone through all the vulnerabilities detected by the scanners and labeled them as:
@@
We have labeled all vulnerabilities as duplicate if Bearer already detected the vulnerability and have not included them in the evaluation.
@@
TABLE III: Comparison of Measures
The results in Table 3 show that passing the Bearer findings to the LLM significantly increases the capabilities of LLM’s in finding vulnerabilities that have not been found by a static scanner. Meanwhile, the Combined LSAST-Approach is not as high-performing as the raw LSAST-Approach, and the raw LLM performs worst.
Interpretation
Two major takeaways can be drawn from the results of our evaluation:
  1. 1.
    Passing the SAST scanner results to the LLM increases the performance of the LLM in finding new vulnerabilities that can’t be found by the SAST scanners.
  2. 2.
    A knowledge retrieval structure for delivering similar vulnerabilities to the LLM can be counterproductive when not carefully crafted.
Explaination
  1. 1.
    Passing the SAST scanner results to the LLM increases the performance of the LLM in finding new vulnerabilities that can’t be found by the SAST scanners.
Our work has proven a very important learning for using LLM’s in static vulnerability code analysis. By informing the LLM of the vulnerabilities that are already present in the code, the LLM can focus on the analysis of other code segments and/or other vulnerabilities that may be present in the code.
  1. 2
    A knowledge retrieval structure for delivering similar vulnerabilities to the LLM can be counterproductive when not carefully crafted.
Our results have shown that using a RAG infrastructure for knowledge retrieval can be counterproductive for the LLM’s vulnerability scanning capabilities. The vulnerability reports that are being passed to the LLM to guide it in the code analysis and deliver up-to-date vulnerabilities have to be highly relevant to the code segment that needs to be analyzed, or it can distract the LLM’s focus and decrease its performance.
@@ V-C Real-world test
To demonstrate that LSAST is capable of identifying vulnerabilities missed by conventional SAST scanners in real-world scenarios, we scanned the Python package pyDust [8] for vulnerabilities and examined the findings in collaboration with the repository’s authors. PyDust is a complex Python package that implements an entity-attribute-value data model, providing mechanisms for creating, manipulating, and serializing entities with dynamic attributes, along with features for change tracking and artifact management.
After scanning the repository, LSAST correctly identified 2 High-Risk, 1 Medium-Risk, and 3 Low-Risk vulnerabilities that conventional scanners failed to detect. These vulnerabilities could lead to serious security issues, such as code injection or remote code execution in the application if left unfixed.
In addition to the 5 true positives, the scanner also reported 2 false positives, highlighting the importance of double-checking scanner results.
Given that PyDust is a sophisticated piece of software and LSAST was still able to find vulnerabilities, we have demonstrated that LSAST can perform effectively in real-world environments.
@@ VI Conclusion
VI-A Future Work
To improve the knowledge retrieval system, multiple steps must be taken to bring the performance to a state where the LLM benefits from the vulnerability reports without getting distracted:
  • A new and improved source of vulnerabilities should be introduced. This source should include the vulnerable code, the CWE-ID, and a description of the vulnerability. Our current source, HackerOne, had the problem of mixing code and description, which led us to use LLM’s to extract the code from the description. This resulted in low-quality vulnerable code snippets.
  • For the abstraction approach, we have used LLM’s to transform the code into a code abstraction. Instead, algorithms should be used to ensure consistency.
Besides improvements in the knowledge retrieval system, other enhancements and research areas in LSAST can be explored:
  • Chain of thought reasoning should be applied.
  • Research should be conducted on how to scan an entire repository without needing to chunk the request and while preserving context, to extend the LLM’s understanding of the whole repository.
  • Consistency of results should be investigated. One idea would be to apply ensemble methods to aggregate predictions from multiple scans with different hyperparameters, to find reliable answers, minimize false positives, and increase consistency.
VI-B Impact of the paper
Our work has made the following contributions:
  • First, we have significantly enhanced LLM’s capabilities in scanning for vulnerabilities by passing static code scanning findings to an LLM, enabling it to find vulnerabilities that would not have been detected if relying on conservative methods alone. By doing this, we have demonstrated the synergy that can be unlocked when combining these technologies.
  • Second, we have addressed LLM’s limitation in gathering up-to-date vulnerability information and have improved current methods for addressing this issue by introducing new approaches for knowledge retrieval systems. Our combined and abstraction approaches outperform the knowledge retrieval systems presented in other papers. These approaches can now be used as new standards for building and improving knowledge retrieval systems for static vulnerability scanning.
  • Third, we have shown that using a knowledge retrieval system can be counterproductive if not perfectly crafted. Reports that are passed to the LLM can distract its focus if they are not highly relevant.
  • Fourth, by using open-source LLM’s in our research, we have mitigated privacy and security risks and have demonstrated that everyone can make use of the capabilities of LLM’s in vulnerability scanning without compromising in these areas.
  • Fifth, we have demonstrated in a real-world scenario that LSAST scanners can detect vulnerabilities missed by traditional scanners, making them essential components of any thorough security scanning process moving forward.
Overall, we have made important findings in the area of LLM-supported static code analysis. We have demonstrated the potential for synergy between traditional static analysis tools and LLM’s and have shown how to utilize such synergy.
While there is still much research to be done in this field, we can confidently say that LLM’s can play a crucial part of the static code analysis process, and we expect to see significant improvements in the near future.

### v2

@@ IV-C Preparing Vulnerability Reports for Knowledge Retrieval
To determine the most relevant reports for the target code, we introduce LLM4Vuln [24], which performs a similarity search based on code functionality. The process begins by summarizing the functionality of each code snippet in the vulnerability reports and storing these summaries in a vector database using the LLM (see Listing 6). When scanning the target code (TC) for vulnerabilities, we also summarize its functionality using the LLM. We then conduct a similarity search in the vector database to find vulnerable code snippets (VC) with similar functionalities. Multiple functionalities of vulnerable code are correlated back to their respective reports using sequence numbers stored in the database, allowing us to extract and deliver the most relevant reports to the LLM.
@@
Recognizing the weak correlation between code functionality and vulnerabilities, we developed a new approach for conducting similarity searches. A stronger correlation often emerges when examining code abstraction in relation to vulnerabilities. Abstraction, or code minimization, simplifies source code to essential components necessary for execution. To perform a similarity search based on abstraction, we extract code from vulnerability reports and apply code minification techniques using minifiers or LLM capabilities (see Listing 7). After minifying the vulnerable code snippets, we store them in a vector database.
@@
When scanning the target code (TC) for vulnerabilities, we first minimize the TC and then search for vulnerable code (VC) with similar minified representations. We aggregate the code abstractions and correlate them back to the original reports using sequence numbers, as previously explained. This approach ensures we gather reports with structures highly resembling the target code, increasing the likelihood of identifying identical vulnerabilities present in both VC and TC.
This method closely aligns with conservative SAST scanner methodologies, which aim to identify vulnerable patterns within code by leveraging predefined rules and heuristics to detect known weaknesses, ensuring thorough and systematic analysis of potential security issues.
@@ IV-D Knowledge Retrieval Methods
To address the issue of LLMs relying on static training data, we use a dynamic Knowledge Retrieval System.
In this section, we provide some illustrative figures to show the various methods available for constructing the knowledge retrieval process through the combination of the functionality approach, abstraction approach, and conservative SAST scanner results.
All the notations are summarized in the Table I.
TABLE I: Definitions of Terms
Abbreviation Description
TC Target Code
SR SAST-Scanner Result
VR Vulnerability Reports
FVR Functionality-similar Vulnerability Reports
CVR Code-Abstraction-similar Vulnerability Reports
k Value of top $k$ in the similarity search
The following approaches are designed to progressively improve the LLM’s ability to identify vulnerabilities by integrating various sources of information:
IV-D1 Raw Own-Knowledge Approach (Baseline)
Figure 1: RAW Own-Knowledge Approach.
@@ IV-D1 Raw Own-Knowledge Approach (Baseline)
This simplest method involves directly querying the LLM using only the TC without any support from external sources or retrieval systems, as shown in Fig. 1. It relies solely on static training data, often missing newer or context-specific vulnerabilities. This approach is considered as the baseline.
IV-D2 RAW Own-Knowledge LSAST-Approach
Figure 2: RAW Own-Knowledge LSAST-Approach.
@@ IV-D2 RAW Own-Knowledge LSAST-Approach
Building upon the baseline, we incorporate the SAST scanner result, SR, along with the TC into the LLM input as our raw LSAST approach, as shown in Fig. 2. This allows the LLM gain initial insights into known vulnerabilities identified by traditional static analysis to enhance the overall detection accuracy.
IV-D3 Functionality-Based Retrieval Approach
Figure 3: RAW Functionality Approach.
@@ IV-D3 Functionality-Based Retrieval Approach
In this approach, depicted in Fig. 3, the LLM first summarizes the functionality of the TC. A similarity search is then conducted within a vector database containing functionalities of known vulnerable code snippets. The most relevant FVR are retrieved and provided to the LLM along with the target code.
IV-D4 Functionality LSAST-Approach
Figure 4: Functionality LSAST-Approach.
@@ IV-D4 Functionality LSAST-Approach
For the functionality, our approach utilizes Eq. (2) to extend the retrieval process by also incorporating SAST results (SR) alongside functionality reports (FVR) as shown in Fig. 4. Combining the detailed vulnerability reports with SAST’s known patterns provides a dual-layered context, improving the LLM’s performance by cross-referencing functionality-driven insights with existing static analysis results.
IV-D5 Abstraction-Based Retrieval Approach
Figure 5: RAW Abstraction Approach
@@ IV-D5 Abstraction-Based Retrieval Approach
In the abstraction approach as shown in Fig.5, the LLM abstracts the TC into its fundamental structual components. A similarity search is conducted within a vector database of code abstractions, retrieving the closest matching CVR.
IV-D6 Abstraction LSAST-approach
Figure 6: Abstraction LSAST-Approach.
@@ IV-D6 Abstraction LSAST-approach
As shown in Fig. 6, our abstraction LSAST-approach combines the strengths of abstraction-based retrieval with SR, providing the LLM with a comprehensive input that includes both structural patterns and known vulnerabilities. This dual approach allows the LLM to validate structural vulnerabilities against established patterns found by SAST, enhancing detection accuracy, especially in complex or poorly documented code segments.
IV-D7 Combined Approach
Figure 7: RAW Combined Approach.
@@ IV-D7 Combined Approach
This approach combines both FVR and CVR, integrating them with the target code to provide a holistic view, demonstrated in Fig. 7. k VR’s from the Functionality similarity search and k VR’s from the Pattern similarity search are gathered. These vulnerability reports are then provided to the LLM.
Figure 8: Combined LSAST-Approach.
@@
The most comprehensive method, Combined-LSAST, integrates TC with functionality (FVR), abstraction (CVR), and bearer SAST scanner results (SR), illustrated in Fig. 8. This approach unites all previously described strategies, leveraging the strengths of each to provide a robust input that covers all angles of vulnerability detection. By combining these methods, the LLM is equipped with detailed insights that span functionality, structure, and known vulnerabilities, resulting in a significantly enhanced detection capability.
@@ V Evaluation
Since previous work has already shown that LLMs can be effectively utilized for vulnerability scanning and it is very time-consuming to evaluate all approaches, we have decided to structure our approach in the most efficient way. First, for all evaluations, we are using Llama-3-70B, as it is a high-performing mid-sized model that can be hosted without needing very large cloud infrastructures or sending the code to third-party LLM providers. In addition, we have first evaluated which of the RAG approaches is most efficient in a small-scale evaluation before attempting the big evaluation between the raw models and the best-performing RAG approach.
V-A Comparison of Abstraction, Functionality, and Combined
We start by evaluating the Abstraction Approach (Eq. (5)), Functionality Approach (Eq. (3)), and the Combined Approach (Eq. (7)) without feeding the bearer input into the prompts given to LLM. We evaluate our approaches using four well-known open-source security testing projects that are described in the following.
  1. 1.
    DVWA (Damn Vulnerable Web Application) [^2] https://github.com/digininja/DVWA: A PHP/MySQL web application intentionally is designed with vulnerabilities to help security professionals test their tools and hone their vulnerability detection skills.
  2. 2.
    DVNA (Damn Vulnerable NodeJS Application) [^3] https://github.com/appsecco/dvna: A Node.js web application that intentionally incorporates OWASP Top 10 Vulnerabilities to demonstrate how these vulnerabilities can be identified and mitigated.
  3. 3.
    OWASP Juice Shop [^4] https://owasp.org/www-project-juice-shop/: A modern JavaScript-based application (using Node.js, Express, and Angular) intentionally built with vulnerabilities. It serves as a platform for security training and testing of security tools.
  4. 4.
    WebGoat  [^5] https://owasp.org/www-project-webgoat/: A Java-based web application deliberately including vulnerabilities, designed for security professionals to explore and understand common Java-specific security issues.
The projects include different programming languages like Java, JavaScript, TypeScript, and PHP, and are built to include a wide range of different vulnerabilities. Across all 4 projects, using our 3 scanner approaches, we have scanned a total of 10 files that include more than 15 different kinds of vulnerabilities.
TABLE II: Comparison of Abstraction, Functionality, and Combined Approaches
@@ V-A Comparison of Abstraction, Functionality, and Combined
For our evaluation metrics, we use standard measures in vulnerability detection:
@@
We use the Bearer results as the reference standard, challenging the three approaches to identify as many vulnerabilities as the static scanner detects. This evaluation aims to determine which of the three approaches best covers the underlying vulnerabilities.
Table II illustrates that the Combined Approach has outperformed both the Abstraction Approach and the Functionality Approach. The 100% FP-Rate is achieved since all files include at least 1 vulnerability, and therefore a scan never had the opportunity to label a file as false positive.
As explained earlier, there is a higher correlation between code abstraction and underlying vulnerabilities compared to functionality and underlying vulnerabilities. The reason is that functionalities can differ greatly but still have the same vulnerabilities, and vice versa. Conservative scanners also scan for underlying code patterns or syntactic features. Therefore, we can see that our new Abstraction and Combined approaches work better than the Functionality approach, which was introduced in LLM4Vuln [24].
@@
TABLE III: Comparison of similarity search for Abstraction and Functionality Approaches
@@
Table III represents the results of the Similarity Search Performance. The TP-Rate can be interpreted as how many of the vulnerabilities that we passed to the LLM were also found by Bearer. In other words, around 67% of the reports that we passed to the LLM, with the abstraction approach, were related to other vulnerabilities than the ones found by Bearer. This indicates that we effectively distracted the LLM from identifying the vulnerabilities detected by Bearer by providing a large number of unrelated vulnerabilities, highlighting the importance of refining the input to ensure the LLM remains focused on the most relevant security threats.
@@ V-B RAW vs LSAST vs LSAST-Combined
We have evaluated that the combined approach performs the best out of the three knowledge retrieval approaches in the above section. To further explore its effectiveness, we compare the combined approach with those that do not incorporate a knowledge retrieval system. We have scanned the open-source projects DVWA, DVNA, Juice Shop, and WebGoat, using three approaches: the combined LSAST-Approach (Eq.(8)) (with knowledge retrieval and bearer results), the raw LSAST-Approach (Eq.(2)) (without knowledge retrieval, but with bearer results), and the raw Approach (Eq.(1)) (without knowledge retrieval and without bearer results). In the combined and RAW LSAST approaches, the Bearer results are directly provided to the LLM. We aimed to determine how many true vulnerabilities the LLM could identify beyond those already detected by Bearer. To classify the vulnerabilities found by the scanners, we manually review each detected vulnerability and labeled them as follows:
@@
Vulnerabilities already detected by Bearer are marked as duplicates and excluded from the evaluation.
TABLE IV: Comparison of Measures.
@@
The results in Table IV indicate that including Bearer findings significantly enhances the LLM’s ability to detect vulnerabilities that static scanners miss. However, the RAW LSAST approach outperforms the Combined LSAST approach, while the RAW approach yields the lowest performance.
Two key insights emerge from our evaluation: first, providing SAST scanner results to the LLM enhances its capability to identify new vulnerabilities that cannot be detected by SAST scanners alone. Second, a knowledge retrieval framework that delivers similar vulnerabilities to the LLM can negatively impact performance if not carefully configured.
Integrating SAST scanner results help the LLM focus on detecting vulnerabilities in code areas not previously identified by Bearer, thereby improving overall detection accuracy. Conversely, our findings indicate that using a RAG infrastructure can be counterproductive. If the vulnerability reports provided to the LLM lack high relevance to the analyzed code segments, they may divert the LLM’s attention and reduce its effectiveness.
@@ V-C Real-world test
To demonstrate that LSAST is capable of identifying vulnerabilities missed by conventional SAST scanners in real-world scenarios, we scan the Python package pyDust  [^6] PyDust (https://github.com/MondoAurora/pydust/tree/main/dust) is a complex Python package that implements an entity-attribute-value data model, providing mechanisms for creating, manipulating, and serializing entities with dynamic attributes, along with features for change tracking and artifact management. for vulnerabilities and examine the findings in collaboration with the repository’s authors.
After scanning the repository, LSAST correctly identifies 2 High-Risk, 1 Medium-Risk, and 3 Low-Risk vulnerabilities that conventional scanners failed to detect. These vulnerabilities could lead to serious security issues, such as code injection or remote code execution in the application if left unfixed.
In addition to the 5 true positives, the scanner also reports 2 false positives, highlighting the importance of double-checking scanner results.
Given that PyDust is a sophisticated piece of software and LSAST is still able to find vulnerabilities, we have demonstrated that LSAST can perform effectively in real-world environments.
@@ VI Conclusion
This study introduces LSAST, integrating LLMs with static application security testing to enhance vulnerability detection. By combining SAST results and a knowledge retrieval system, LSAST improves the detection of complex vulnerabilities beyond the capabilities of traditional scanners.
Our results show that integrating SAST with LLMs significantly boosts detection accuracy, though the performance of knowledge retrieval depends heavily on the relevance of the data provided. The research addresses key challenges, including privacy concerns and the limitations of static training data, contributing to more effective cybersecurity analysis. Moreover, the real-world test on the Python package pyDust validates the practical applicability of LSAST, showing its capability to identify high-risk vulnerabilities that conventional scanners miss.
Future work will refine knowledge retrieval strategies and explore deeper integration of LLMs with dynamic testing methods to further strengthen vulnerability detection systems. Another promising direction for future research is the development of adaptive learning mechanisms that allow LLMs to continuously update their knowledge base with the latest security threats and mitigation techniques.